Calculate &Go at 125 C for the following reaction. Assume that ΔΗο and temperature independent. 5. ΔΟ are 2CH,OH(g) +302(9) 2C02(g) +4H200) Compound AS H2O CHOH 200 70 393.5 -286 201 240 205
Answer #1

?Hºrxn = 4?HºH2O + 2?HºCO2 - 3?HºO2 - 2?HºCH3OH

=(4x -286) + (2x-393.5) - 0 - (2x-201)

=-1529 kJ/mol = -1529000 J /mol

?Sºrxn = 4?SºH2O + 2?SºCO2 - 3?SºO2 - 2?SºCH3OH

=(4x70) + (2x214) -(3x205) - (2x240)

= -387 J/K/mol

T = 125+273.15 = 398.15 K

?Gº = ?Hº-T?Sº = -1529000 - ( 398.15 x -387 )

= 1374916 J / mol

= 1374.9 kJ /mol